OpenStudy (anonymous):

write a system of equations with the solution (4, -3)

OpenStudy (whpalmer4):

That's pretty easy. Just make up some equations with variables \(x,y\) that use that solution. For example, say we needed a solution of (1,2): Some randomly chosen equations might be x + y = ??? 3x + 4y = ??? Plug in x=1, y = 2 and figure out the values to replace the ??? 1+2 = 3 3(1) + 4(2) = 11 so one system of equations with solution (1,2) is \[x+y = 3\]\[3x+4y=11\]